# Greedy Low-Rank Gradient Compression for Distributed Learning with Convergence Guarantees

### v2

@@ Abstract
Index Terms:
Distributed Learning, Low-Rank Compression, Communication-Efficient Optimization, Error Feedback.
@@ III Preliminary
Notation. Let $\mathds{1}_{n}$ denote the vector of all-ones of $n$ dimensions and $\bm{I}_{n}\in\mathbb{R}^{n\times n}$ the identity matrix. We introduce the set $[n]:=\{1,\cdots,n\}$. Given the index set $\mathcal{J}$, notation $\bm{U}_{:,\mathcal{J}}$ refers to the submatrix of $\bm{U}$ formed by selecting the columns of $\bm{U}$ indexed by the set $\mathcal{J}$. The notation $\bm{U}_{:,:r}$ refers to the first $r$ columns of the matrix $\bm{U}$.
@@ V-A Error feedback
$\bm{R}^{(i)}_{t}=\bm{P}^{\top}_{t}(\bm{G}_{t}^{(i)}+\bm{E}_{t-1}^{(i)}),\bm{R}_{t}=\frac{1}{N}\sum_{i=1}^{N}\bm{R}^{(i)}_{t},\hat{\bm{G}}_{t}=\bm{P}_{t}\bm{R}_{t},$ (10)
@@
Input: $N$ nodes, learning rate $\gamma$, number of total iterations $T$, subspace changing frequency $\tau$, rank $r$, error buffer $\bm{E}_{-1}=\bm{0}$, weight $\bm{X}_{0}\in\mathbb{R}^{m\times n}$ and projection matrix $\bm{P}_{-1}\in\mathbb{R}^{m\times r}$ on each node $i\in[N]$ with shape $m\leq n$.
Output: Sequence of model weights $\{\bm{X}_{t}\}_{t=0}^{T+1}$.
@@ V-B Semi-lazy SVD update
=$\overset{(a)}{=}$ $\sum\limits_{j\notin\mathcal{J}}\|\bm{u}_{j}^{\top}\bm{G}_{t}-\bm{0}^{\top}\|_{F}^{2}+\sum\limits_{j\in\mathcal{J}}\|\bm{u}_{j}^{\top}\bm{G}_{t}-\bm{u}_{j}^{\top}\bm{G}_{t}\|_{F}^{2}$
@@
Input: $N$ nodes, rank $r$, global orthogonal matrix $\bm{U}\in\mathbb{R}^{m\times m}$, local gradient $\bm{G}_{t}^{(i)}\in\mathbb{R}^{m\times n}$ for $i\in[N]$
@@
Compute $\overline{\bm{\Lambda}}\leftarrow\frac{1}{N}\sum_{i=1}^{N}{\bm{\Lambda}}^{(i)}$. (All-Reduce)
Compute $\overline{\bm{\Sigma}}\leftarrow(\overline{\bm{\Lambda}})^{2}$ with element-wise square operation.
@@ V-C Approximate global Top-$r$ projection
Input: $N$ nodes, number of total iterations $T$, subspace changing frequency $\tau$, rank $r$, $\bm{X}_{0}\in\mathbb{R}^{m\times n}$ and $\bm{E}_{-1}^{(i)}=\bm{0}$ for node $i\in[N]$ with shape $m\leq n$. Initialize $\bm{U}=\bm{I}_{m}$.
Output: Sequence of model weights $\{\bm{X}_{t}\}_{t=0}^{T+1}$.
@@
[…]
@@ VI Convergence Analysis
Under Assumptions 1, 2, 4 and 5, if we let $\gamma\leq c_{l}/(4Lc_{u}^{2})$, GreedyLore with Adam converges as $\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}[\|\nabla f(\bm{X}_{t})\|_{F}^{2}]\leq\frac{2\Delta_{0}}{\gamma c_{l}(T+1)}+\frac{40\gamma^{2}L^{2}c_{u}^{4}B_{s}^{2}}{(1-\beta_{1})^{2}c_{l}^{2}\delta}$ $+\frac{4c_{u}(2+\gamma Lc_{u})\tau B_{s}^{2}}{c_{l}(1-\beta_{1})(T+1)}+\frac{4\gamma L\tau^{2}c_{u}^{2}B_{s}^{2}}{(1-\beta_{1})^{2}c_{l}(T+1)}+\frac{2\gamma Lc_{u}^{2}\sigma^{2}}{Nc_{l}},$ where $\delta:=r/m$, $\Delta_{0}:=f(\bm{X}_{0})-\inf_{\bm{X}}f(\bm{X})$. If we further choose $\gamma$ properly, GreedyLore with Adam converges as $\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}[\|\nabla f(\bm{X}_{t})\|_{F}^{2}]=$ (32) $\mathcal{O}\left(\sqrt{\frac{L\Delta_{0}\sigma^{2}}{N(T+1)}}+\sqrt[3]{\frac{L^{2}\Delta_{0}^{2}B_{s}^{2}}{\delta(T+1)^{2}}}+\frac{\tau B_{s}(B_{s}+\sqrt{\Delta_{0}})+L\Delta_{0}}{T+1}\right).$
@@ VIII Conclusion
In this work, we presents GreedyLore, a communication-efficient, greedy low-rank gradient compression algorithm designed to address fundamental limitations in existing distributed optimization methods. By combining a novel greedy projector selection strategy with an integrated error-feedback mechanism, GreedyLore achieves strong theoretical convergence guarantees without imposing restrictive assumptions such as bounded gradients. Our rigorous analysis confirms a convergence rate consistent with standard adaptive optimization methods. Empirical evaluations across multiple benchmark tasks, including ResNet pre-training, LLaMA model pre-training and RoBERTa fine-tuning, validate GreedyLore’s superior performance relative to existing compression methods. Moreover, GreedyLore maintains negligible additional memory overhead and integrates with established distributed training frameworks, highlighting its practical utility.
@@ Appendix A Proofs for convergence theorems
[…]
@@
$+\frac{\gamma^{2}L}{2}\mathbb{E}\left[\left\|\bm{\Gamma}_{t}\odot\widetilde{\bm{G}}_{t}-\Delta\bm{\Gamma}_{t}\odot\left(\frac{\beta_{1}}{1-\beta_{1}}\bm{M}_{t-1}+\overline{\bm{E}}_{t-1}\right)\right\|_{F}^{2}\right]$
@@
$+\frac{\gamma^{3}L^{2}c_{u}^{4}}{c_{l}}\mathbb{E}\left[\left\|\frac{\beta_{1}}{1-\beta_{1}}\bm{M}_{t-1}+\overline{\bm{E}}_{t-1}\right\|_{F}^{2}\right]+\frac{\gamma^{2}Lc_{u}^{2}\sigma^{2}}{n},$ (A8)
@@
$+\gamma L\mathbb{E}\left[\left\|\bm{\Gamma}_{t}\odot\left(\frac{\beta_{1}}{1-\beta_{1}}\bm{M}_{t-1}+\overline{\bm{E}}_{t-1}\right)\right\|_{F}\|\Delta\bm{\Gamma}_{t}\odot\widetilde{\bm{G}}_{t}\|_{F}\right]$
@@
$\leq$ $\left(B_{s}^{2}+\frac{\tau B_{s}^{2}}{1-\beta_{1}}+\frac{\gamma Lc_{u}\tau B_{s}^{2}}{1-\beta_{1}}\right)\cdot\frac{2}{c_{l}(T+1)}\sum_{t=0}^{T}\mathbb{E}[\|\Delta\bm{\Gamma}_{t}\|_{\max}]$
$+\frac{2[f(\bm{X}_{0})-\mathbb{E}[f(\bm{Y}_{T+1})]]}{\gamma c_{l}(T+1)}+\left(\frac{12\beta_{1}^{2}}{(1-\beta_{1})^{2}}+8\right)\frac{2\gamma^{2}L^{2}c_{u}^{4}B_{s}^{2}}{c_{l}^{2}\delta}$
@@
Following the proof of (A8) in Theorem 2, while substituting $\Delta\bm{\Gamma}_{t}$ with $\bm{0}$, $\bm{\Gamma}_{t}$ with $\bm{1}$, $c_{u}$ with $1$, $c_{l}$ with 1, and Lemma 4 with Lemma 3, we obtain
@@
$+\left(\frac{12\beta_{1}^{2}}{(1-\beta_{1})^{2}}+8\right)\frac{\gamma^{3}L^{2}(B^{2}+\sigma^{2})}{\delta},$ (A9)
@@ Appendix B Missing proofs
WLOG assume $i=1$, $j=2$ and $\sigma_{1}\geq\sigma_{2}$. It suffices to prove that for any given ${\bm{\xi}}_{3},\cdots,{\bm{\xi}}_{m}\in\mathbb{R}$, the conditional probabilities satisfy $\mathbb{P}_{A}[\Lambda_{1}]\geq\mathbb{P}_{A}[\Lambda_{2}]$, where $\mathbb{P}_{A}[\cdot]:=\mathbb{P}[\cdot|{\bm{\xi}}_{3},\cdots,{\bm{\xi}}_{m}]$, and $\Lambda_{i}:=\{|{\bm{\xi}}_{i}|\in\mathrm{Top}_{k}(|{\bm{\xi}}_{1}|,\cdots,|{\bm{\xi}}_{m}|)\}$. WLOG assume $|{\bm{\xi}}_{3}|\leq|{\bm{\xi}}_{4}|\leq\cdots\leq|{\bm{\xi}}_{m}|$. If $k=m$ or $\sigma_{2}=0$, the result is trivial. In the following we assume $k<m$ and $\sigma_{2}>0$. If $k=m-1$, it holds that
@@
which implies $\mathbb{P}_{A}[\Lambda_{1}]\geq\mathbb{P}_{A}[\Lambda_{2}]$. If $22\leq k\leq m-$ 2, it holds that
@@
$=$ $\mathbb{E}[\|\nabla f(\bm{X}_{t})+\overline{\bm{E}}_{t-1}\|_{F}^{2}]+\mathbb{E}\left[\left\|\frac{1}{N}\sum_{i=1}^{N}\nabla F_{i}(\bm{X}_{t};{\bm{\xi}}_{t}^{(i)})-\nabla f(\bm{X}_{t})\right\|_{F}^{2}\right]$
@@
(${\leq}$ $\frac{4-2\delta-2\delta^{2}}{\delta^{2}}B^{2}+\frac{2(1-\delta)\sigma^{2}}{N\delta},$ (A16)
@@
$\|\Delta\bm{\Gamma}_{t}\|_{\max}=$ $\|\bm{\Gamma}_{t-1}-\bm{\Gamma}_{t}\|_{\max}\leq\max_{1\leq i\leq m\atop 1\leq j\leq n}\{{{\Gamma}}_{t-1,i,j}\}-\min_{1\leq i\leq m\atop 1\leq j\leq n}\{{{\Gamma}}_{t,i,j}\}$

### v4

@@ III Preliminary
Notation. Let $\mathds{1}_{n}$ denote the vector of all-ones of $n$ dimensions and $\bm{I}_{n}\in\mathbb{R}^{n\times n}$ the identity matrix. We introduce the set $n]:=\{1,\cdots,n\}$. Given the index set $\mathcal{J}$, notation $\bm{U}_{:,\mathcal{J}}$ refers to the submatrix of $\bm{U}$ formed by selecting the columns of $\bm{U}$ indexed by the set $\mathcal{J}$. The notation $\bm{U}_{:,:r}$ refers to the first $r$ columns of the matrix $\bm{U}$.
@@ V-A Error feedback
$\hskip-2.84526pt\bm{R}^{(i)}_{t}\hskip-1.42262pt=\hskip-1.42262pt\bm{P}^{\top}_{t}(\bm{G}_{t}^{(i)}\hskip-1.42262pt+\hskip-1.42262pt\bm{E}_{t\hskip-1.42262pt-\hskip-1.42262pt1}^{(i)}),\bm{R}_{t}\hskip-1.42262pt=\hskip-1.42262pt\frac{1}{N}\sum_{i=1}^{N}\bm{R}^{(i)}_{t},\hat{\bm{G}}_{t}\hskip-1.42262pt=\hskip-1.42262pt\bm{P}_{t}\bm{R}_{t},$ (10)
@@
Input: NN nodes, learning rate $\gamma$, number of total iterations $T$, subspace changing frequency $\tau$, rank $r$, error buffer $\bm{E}_{-1}=\bm{0}$, weight $\bm{X}_{0}\in\mathbb{R}^{m\times n}$ and projection matrix $\bm{P}_{-1}\in\mathbb{R}^{m\times r}$ on each node $i\in[N]$ with shape $m\leq n$.
2pt Output: Sequence of model weights $\{\bm{X}_{t}\}_{t=0}^{T+1}$.
@@ V-B Semi-lazy SVD update
$\overset{(a)}{=}$ $\sum\limits_{j\notin\mathcal{J}}\|\bm{u}_{j}^{\top}\bm{G}_{t}-\bm{0}^{\top}\|_{F}^{2}+\sum\limits_{j\in\mathcal{J}}\|\bm{u}_{j}^{\top}\bm{G}_{t}-\bm{u}_{j}^{\top}\bm{G}_{t}\|_{F}^{2}$
@@
Input: NN nodes, rank $r$, global orthogonal matrix $\bm{U}\in\mathbb{R}^{m\times m}$, local gradient $\bm{G}_{t}^{(i)}\in\mathbb{R}^{m\times n}$ for $i\in[N]$
@@
0.5mm Compute $\overline{\bm{\Lambda}}\leftarrow\frac{1}{N}\sum_{i=1}^{N}{\bm{\Lambda}}^{(i)}$. (All-Reduce)
0.5mm Compute $\overline{\bm{\Sigma}}\leftarrow(\overline{\bm{\Lambda}})^{2}$ with element-wise square operation.
@@ V-C Approximate global Top-$r$ projection
Input: NN nodes, number of total iterations $T$, subspace changing frequency $\tau$, rank $r$, $\bm{X}_{0}\in\mathbb{R}^{m\times n}$ and $\bm{E}_{-1}^{(i)}=\bm{0}$ for node $i\in[N]$ with shape $m\leq n$. Initialize $\bm{U}=\bm{I}_{m}$.
2pt Output: Sequence of model weights $\{\bm{X}_{t}\}_{t=0}^{T+1}$.
@@
 $\bm{E}_{t}^{(i)}\leftarrow\bm{G}_{t}^{(i)}\hskip-2.84526pt-\hskip-2.84526pt\bm{P}_{t}\bm{R}_{t}^{\top}\ \mathsf{\bf if\ }\mathsf{mod}(t,\tau)\neq 0\ \mathsf{\bf otherwise}\ \bm{0}$.
@@ VI Convergence Analysis
Under Assumptions 1, 2, 4 and 5, if we let $\gamma\leq c_{l}/(4Lc_{u}^{2})$, GreedyLore with Adam converges as $\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}[\|\nabla f(\bm{X}_{t})\|_{F}^{2}]\leq\frac{2\Delta_{0}}{\gamma c_{l}(T+1)}+\frac{40\gamma^{2}L^{2}c_{u}^{4}B_{s}^{2}}{(1-\beta_{1})^{2}c_{l}^{2}\delta}$ $+\frac{4c_{u}(2+\gamma Lc_{u})\tau B_{s}^{2}}{c_{l}(1-\beta_{1})(T+1)}+\frac{4\gamma L\tau^{2}c_{u}^{2}B_{s}^{2}}{(1-\beta_{1})^{2}c_{l}(T+1)}+\frac{2\gamma Lc_{u}^{2}\sigma^{2}}{Nc_{l}},$ where $\delta:=r/m$, $\Delta_{0}:=f(\bm{X}_{0})-\inf_{\bm{X}}f(\bm{X})$. If we further choose $\gamma$ properly, GreedyLore with Adam converges as $\frac{1}{T+1}\hskip-1.42262pt\sum_{t=0}^{T}\mathbb{E}[\|\nabla f(\bm{X}_{t})\|_{F}^{2}]=$ (32) $\mathcal{O}\hskip-1.70717pt\left(\hskip-1.70717pt\sqrt{\frac{L\Delta_{0}\sigma^{2}}{N(T+1)}}\hskip-1.70717pt+\hskip-1.70717pt\sqrt[3]{\frac{L^{2}\Delta_{0}^{2}B_{s}^{2}}{\delta(T+1)^{2}}}\hskip-1.70717pt+\hskip-1.70717pt\frac{\tau B_{s}(B_{s}\hskip-1.70717pt+\hskip-1.70717pt\sqrt{\Delta_{0}})\hskip-1.70717pt+\hskip-1.70717ptL\Delta_{0}}{T+1}\hskip-1.70717pt\right).\hskip-4.2679pt$
@@ VIII Conclusion
In this work, we presents GreedyLore, a communication-efficient, greedy low-rank gradient compression algorithm designed to address fundamental limitations in existing distributed optimization methods. By combining a novel greedy projector selection strategy with an integrated error-feedback mechanism, GreedyLore achieves strong theoretical convergence guarantees without imposing restrictive assumptions. Our rigorous analysis confirms a convergence rate consistent with standard adaptive optimization methods. Empirical evaluations across multiple benchmark tasks, including ResNet pre-training, LLaMA model pre-training and RoBERTa fine-tuning, validate GreedyLore’s superior performance relative to existing compression methods. Moreover, GreedyLore maintains negligible additional memory overhead and integrates with established distributed training frameworks, highlighting its practical utility.
@@ Appendix A Proofs for convergence theorems
$\ \leq$ $\hskip-1.99168pt-\hskip-1.99168pt\gamma\mathbb{E}[\hskip-1.42262pt\langle\nabla\hskip-1.42262ptf(\hskip-1.42262pt\bm{X}_{t}\hskip-1.42262pt),\hskip-1.42262pt\bm{\Gamma}\hskip-1.42262pt_{t}\hskip-1.42262pt\odot\hskip-1.42262pt\widetilde{\bm{G}}_{t}\rangle\hskip-1.42262pt]\hskip-1.99168pt-\hskip-1.99168pt\gamma\mathbb{E}[\hskip-1.42262pt\langle\nabla\hskip-1.42262ptf(\hskip-1.42262pt\bm{Y}_{t}\hskip-1.42262pt)\hskip-1.99168pt-\hskip-1.99168pt\nabla\hskip-1.42262ptf(\hskip-1.42262pt\bm{X}_{t}\hskip-1.42262pt),\hskip-1.42262pt\bm{\Gamma}\hskip-1.42262pt_{t}\hskip-1.42262pt\odot\hskip-1.42262pt\widetilde{\bm{G}}_{t}\rangle\hskip-1.42262pt]$
@@
$+\frac{\gamma^{2}L}{2}\mathbb{E}\hskip-1.42262pt\left[\hskip-1.42262pt\left\|\bm{\Gamma}_{t}\odot\widetilde{\bm{G}}_{t}\hskip-1.42262pt-\hskip-1.42262pt\Delta\bm{\Gamma}_{t}\odot\left(\hskip-1.42262pt\frac{\beta_{1}}{1\hskip-1.42262pt-\hskip-1.42262pt\beta_{1}}\bm{M}_{t\hskip-1.42262pt-\hskip-1.42262pt1}\hskip-1.42262pt+\hskip-1.42262pt\overline{\bm{E}}_{t\hskip-1.42262pt-\hskip-1.42262pt1}\hskip-1.42262pt\right)\hskip-1.42262pt\right\|_{F}^{2}\hskip-1.42262pt\right]$
@@
$+\frac{\gamma^{3}L^{2}c_{u}^{4}}{c_{l}}\mathbb{E}\hskip-1.42262pt\left[\hskip-1.42262pt\left\|\hskip-1.42262pt\frac{\beta_{1}}{1\hskip-1.42262pt-\hskip-1.42262pt\beta_{1}}\bm{M}_{t\hskip-1.42262pt-\hskip-1.42262pt1}\hskip-1.42262pt+\hskip-1.42262pt\overline{\bm{E}}_{t\hskip-1.42262pt-\hskip-1.42262pt1}\hskip-1.42262pt\right\|_{F}^{2}\hskip-1.42262pt\right]\hskip-1.42262pt+\hskip-1.42262pt\frac{\gamma^{2}Lc_{u}^{2}\sigma^{2}}{n},$ (A8)
@@
$+\gamma L\mathbb{E}\left[\left\|\bm{\Gamma}_{t}\hskip-1.42262pt\odot\hskip-1.42262pt\left(\frac{\beta_{1}}{1-\beta_{1}}\bm{M}_{t-1}\hskip-1.42262pt+\hskip-1.42262pt\overline{\bm{E}}_{t-1}\right)\right\|_{F}\|\Delta\bm{\Gamma}_{t}\odot\widetilde{\bm{G}}_{t}\|_{F}\right]$
@@
$\leq$ $\left(B_{s}^{2}\hskip-1.42262pt+\hskip-1.42262pt\frac{\tau B_{s}^{2}}{1\hskip-1.42262pt-\hskip-1.42262pt\beta_{1}}\hskip-1.42262pt+\hskip-1.42262pt\frac{\gamma Lc_{u}\tau B_{s}^{2}}{1\hskip-1.42262pt-\hskip-1.42262pt\beta_{1}}\right)\hskip-1.42262pt\cdot\hskip-1.42262pt\frac{2}{c_{l}(T\hskip-1.42262pt+\hskip-1.42262pt1)}\sum_{t=0}^{T}\mathbb{E}[\|\Delta\bm{\Gamma}_{t}\|_{\max}]$
$+\frac{2[f(\bm{X}_{0})\hskip-1.42262pt-\hskip-1.42262pt\mathbb{E}[f(\bm{Y}_{T+1})]]}{\gamma c_{l}(T\hskip-1.42262pt+\hskip-1.42262pt1)}\hskip-1.42262pt+\hskip-1.42262pt\left(\hskip-1.42262pt\frac{12\beta_{1}^{2}}{(1\hskip-1.42262pt-\hskip-1.42262pt\beta_{1})^{2}}\hskip-1.42262pt+\hskip-1.42262pt8\hskip-1.42262pt\right)\hskip-1.42262pt\frac{2\gamma^{2}L^{2}c_{u}^{4}B_{s}^{2}}{c_{l}^{2}\delta}$
@@
Following the proof of (A8) in Theorem 2, while substituting $\Delta\bm{\Gamma}_{t}$ with $\bm{0}$, $\bm{\Gamma}_{t}$ with $\bm{1}$, $c_{u}$ with 11, $c_{l}$ with 1, and Lemma 4 with Lemma 3, we obtain
@@
$\hskip-35.56593pt+\hskip-1.42262pt\left(\hskip-1.42262pt\frac{12\beta_{1}^{2}}{(1\hskip-1.42262pt-\hskip-1.42262pt\beta_{1})^{2}}\hskip-1.42262pt+\hskip-1.42262pt8\hskip-1.42262pt\right)\hskip-1.42262pt\frac{\gamma^{3}L^{2}(B^{2}\hskip-1.42262pt+\hskip-1.42262pt\sigma^{2})}{\delta},$ (A9)
@@ Appendix B Missing proofs
WLOG assume $i=1$, $j=2$ and $\sigma_{1}\geq\sigma_{2}$. It suffices to prove that for any given ${\bm{\xi}}_{3},\cdots,{\bm{\xi}}_{m}\in\mathbb{R}$, the conditional probabilities satisfy $\mathbb{P}_{A}[\Lambda_{1}]\geq\mathbb{P}_{A}[\Lambda_{2}]$, where $\mathbb{P}_{A}[\cdot]:=\mathbb{P}[\cdot|{\bm{\xi}}_{3},\cdots,{\bm{\xi}}_{m}]$, and $\Lambda_{i}:=\{|{\bm{\xi}}_{i}|\in\mathrm{Top}_{k}(|{\bm{\xi}}_{1}|,\cdots,|{\bm{\xi}}_{m}|)\}$. WLOG assume $|{\bm{\xi}}_{3}|\leq|{\bm{\xi}}_{4}|\leq\cdots\leq|{\bm{\xi}}_{m}|$. If $k=m$ or $\sigma_{2}=0$, the result is trivial. In the following we assume $k<m$ and $\sigma_{2}>0$. If k=m1k=m-1, it holds that
@@
which implies $\mathbb{P}_{A}[\Lambda_{1}]\geq\mathbb{P}_{A}[\Lambda_{2}]$. If $2\leq k\leq m-2$, it holds that
@@
$=$ $\mathbb{E}[\|\nabla\hskip-1.42262ptf(\hskip-1.42262pt\bm{X}\hskip-1.42262pt_{t}\hskip-1.42262pt)\hskip-1.42262pt+\hskip-1.42262pt\overline{\bm{E}}\hskip-1.42262pt_{t\hskip-1.42262pt-\hskip-1.42262pt1}\|_{F}^{2}]\hskip-1.42262pt+\hskip-1.42262pt\mathbb{E}\hskip-1.42262pt\left[\hskip-1.42262pt\left\|\frac{1}{N}\hskip-1.42262pt\sum_{i=1}^{N}\hskip-1.42262pt\nabla\hskip-1.42262ptF_{i}(\hskip-1.42262pt\bm{X}\hskip-1.42262pt_{t};\hskip-1.42262pt{\bm{\xi}}_{t}^{(i)}\hskip-1.42262pt)\hskip-1.42262pt-\hskip-2.84526pt\nabla\hskip-1.42262ptf(\hskip-1.42262pt\bm{X}\hskip-1.42262pt_{t}\hskip-1.42262pt)\hskip-1.42262pt\right\|_{F}^{2}\hskip-1.42262pt\right]$
@@
[…]
@@
$\|\Delta\bm{\Gamma}_{t}\|_{\max}\hskip-1.42262pt=$ $\|\bm{\Gamma}_{t\hskip-1.42262pt-\hskip-1.42262pt1}\hskip-1.42262pt-\hskip-1.42262pt\bm{\Gamma}_{t}\|_{\max}\hskip-1.42262pt\leq\hskip-1.99168pt\max_{1\leq i\leq m\atop 1\leq j\leq n}\hskip-1.99168pt\{\hskip-1.42262pt{{\Gamma}}_{t\hskip-1.42262pt-\hskip-1.42262pt1,i,j}\hskip-1.42262pt\}\hskip-1.42262pt-\hskip-4.2679pt\min_{1\leq i\leq m\atop 1\leq j\leq n}\hskip-1.99168pt\{\hskip-1.42262pt{{\Gamma}}_{t,i,j}\hskip-1.42262pt\}$